# LLM-Deliberation: Evaluating LLMs with Interactive Multi-Agent Negotiation Games

### v1

@@ Abstract
There is a growing interest in using Large Language Models (LLMs) as agents to tackle real-world tasks that may require assessing complex situations. Yet, we have a limited understanding of LLMs’ reasoning and decision-making capabilities, partly stemming from a lack of dedicated evaluation benchmarks. As negotiating and compromising are key aspects of our everyday communication and collaboration, we propose using scorable negotiation games as a new evaluation framework for LLMs. We create a testbed of diverse text-based, multi-agent, multi-issue, semantically rich negotiation games, with easily tunable difficulty. To solve the challenge, agents need to have strong arithmetic, inference, exploration, and planning capabilities, while seamlessly integrating them. Via a systematic zero-shot Chain-of-Thought prompting (CoT), we show that agents can negotiate and consistently reach successful deals. We quantify the performance with multiple metrics and observe a large gap between GPT-4 and earlier models. Importantly, we test the generalization to new games and setups. Finally, we show that these games can help evaluate other critical aspects, such as the interaction dynamics between agents in the presence of greedy and adversarial players.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023b) have been primarily trained in an unsupervised way on massive datasets. Despite that, they perform relatively well in setups beyond traditional NLP tasks, such as using tools (Schick et al., 2023; Patil et al., 2023; Lu et al., 2023a; Yao et al., 2023b) or solving reasoning problems (Srivastava et al., 2023; Wei et al., 2022). This motivated their adoption in many real-world applications (OpenAI, 2023a; Microsoft, 2023b; c) and a recent line of work to leverage LLMs as interactive agents (Liu et al., 2023). Given this discrepancy between training paradigms and these new adoptions, we need new evaluation frameworks that help us understand and systematically test models’ capabilities, limitations, and potential misuse.
Toward this goal, we propose negotiation games as a new evaluation benchmark for LLMs. Our choice of using negotiation is motivated by several factors. From a practical view, negotiation plays a central role in our everyday lives (Kramár et al., 2022), from scheduling meetings (Duplex, ), satisfying customers, and agreeing on contracts to high-stake decisions such as authorizing loans or peace mediation. As AI models and LLMs continue to be increasingly relied on as customer service bots (HBR, ), personal (Duplex, ) and negotiation assistants (Icertis, ; Luminance, ), and even autonomous agents (Pactum, ), it is essential to evaluate their performance in such tasks.
@@
Figure 1: Left: 6 parties negotiate over 5 issues ($A$, $B$, …, $E$) with different sub-options ($b_{1}$, $b_{2}$, etc.). The maximum score any party can achieve is 100. Each party has its own secret scores for the sub-options and a minimum threshold for acceptance. The pie charts represent the priority of issues and the most preferred sub-option. Right: A depiction of how parties can compromise to reach a common agreement that increases their collective average score by finding adjustments to their ideal deal. The graph is the result of one of our experiments with GPT-4. Over rounds, the leading agent $p_{1}$ proposes deals in its turn that reduce its own score (while still being above its own minimum threshold) but increase the average collective score of all agents (which $p_{1}$ cannot directly observe).
From a technical view, interactive negotiation games, with a complex interplay between self-interested parties, provide an excellent platform to characterize and probe for LLMs’ capabilities. Agents must assess the value of deals according to their own goals, have a representation of others’ goals, update this representation based on newer observations, plan and adapt their strategies over rounds, weigh different options, and finally find common grounds. These sub-tasks require substantial and non-trivial arithmetic and strategic reasoning under only partial observations. They are also related to commonsense reasoning (Talmor et al., 2019; Sap et al., 2019) and Theory-of-Mind (ToM) evaluation (Sclar et al., 2023; Sap et al., 2022); agents need to rely on the game’s real-world semantics to ascribe goals to others and must differentiate between their goals and those of others.
@@
We first leverage an existing commonly-used scorable role-play negotiation game (Susskind, 1985; Susskind & Corburn, 2000) with multiple parties and different issues under negotiation. A high-level overview is shown in Figure 1. Parties have their own real-world-inspired goals determining their unique and secret scores for issues. They also have a minimum threshold for agreement. The priorities vary between parties, creating a non-zero-sum game with potential for cooperation. To rule out memorization and provide a rich benchmark, we create semantically equivalent games by perturbing parties/issues names, and we use an LLM as a seed to design 3 new diverse games. The scores and thresholds control the set of feasible solutions, providing a way to quantify performance. This helps create a less saturating benchmark by instantiating new games with varying difficulty levels where there is room for further enhancements and evaluating future models.
We systematically evaluate various Chain-of-Thought (CoT) prompting (Wei et al., 2022) strategies and found that GPT-4 performs significantly better than earlier models. We further use these negotiation games to study agents’ interaction in unbalanced adversarial settings, a critical aspect for future autonomous systems with little human oversight and different entity-controlled LLMs. We show that agents’ behavior can be modulated to promote greediness or attack other agents, frequently sabotaging the negotiation and altering other cooperative agents’ behaviors as well.
Our work takes a step toward using complex interactive games to evaluate LLMs negotiation capabilities, the potential for manipulation, and possible future robustification, building on previous work of using text-based games as a practical testbed (Pan et al., 2023; Zhou et al., 2023; Hendrycks et al., 2022). To foster future research, we will make our toolkit of games and code publicly available at https://github.com/S-Abdelnabi/LLM-Deliberation
2 Related Work
We briefly review related work on evaluating LLMs as agents and in negotiation and other games.
LLMs in negotiation and other games. Fu et al. (2023) proposed a simple bargaining game in which two LLMs negotiate a balloon’s price over multiple rounds. The agents receive AI feedback from a critic. Our task is substantially more complex as it involves multi-agent and multi-issue cooperative negotiation with detailed individual preferences and semantically rich simulation. In addition, Akata et al. (2023) evaluated models on several two-player games, such as Prisoner’s Dilemma, and contrasted LLMs’ performance with human strategies. They found that GPT-4’s behavior can be modulated by prompting it to predict the other player’s actions or asking it to be forgiving. We observe similar patterns in our setup; cooperation increases when instructing agents to reason about others’ preferences and accommodate them. Moreover, there are several other text-based games for evaluating LLMs in tasks other than negotiation, such as social decision-making and ethical behavior (Pan et al., 2023) and digital card games and puzzles (Liu et al., 2023).
Strategic reasoning with LLMs. Gandhi et al. (2023b) studied matrix and two-player negotiation games and proposed a prompting strategy that introduces an explicit structure of search (exploring solution space), value assignment (assigning rewards to actions), and belief tracking (beliefs about other players), accompanied by few-shot demonstrations. In contrast, we use zero-shot CoT. The search step is conceptually similar to prompting strategies such as Tree-of-Thoughts (ToT) (Yao et al., 2023a). Notably, we found that GPT-4 performs relatively well without explicit search; it often autonomously iterates over its own suggested deals until its minimum threshold is met. LMs have been used in previous Diplomacy game frameworks to generate free-form dialogue, integrated with separate planning and reinforcement learning algorithms (Bakhtin et al., 2022). Our work is the first to use LLMs solely in a relatively complex negotiation task.
LLMs as interactive agents. Liu et al. (2023) recently presented AgentBench, a benchmark to evaluate LLMs in 8 distinct environments, including web browsing, web shopping, digital card games, etc. Additionally, Park et al. (2023) introduced a simulation in which agents interact based on their personas and relevant retrieved memories, focusing on social behaviors. We focus instead on negotiation and what it entails in terms of interaction and cooperation toward achieving a goal (e.g., reaching an agreement, etc.). It is useful to note that previous work that studied how interactive strategies can emerge from passive learning (Lampinen et al., 2023), or how an LM can infer goals which in turn affects subsequent generations (Andreas, 2022), may be relevant to explain why LLMs can perform well on our task without needing additional adaptation.
@@ 3 Game Description
The game is based on a negotiation role-play exercise (Susskind, 1985; Susskind & Corburn, 2000) that we further adapt by writing our own description. Besides this game (called “base” in the rest), we created new ones by prompting an LLM (we used Bing Chat (Microsoft, 2023a)) to generate games with a similar style of cooperating and competing interests between parties. All games consist of 6 parties, $P=\{p_{1},p_{2},...,p_{6}\}$, and 5 issues $I=\{A,B,...,E\}$, with the following dynamics.
@@
Parites. An entity $p_{1}$ proposes a project (e.g., an airport, a solar power plant, a new sports park, etc.) that it will manage and invest in and wants to increase the return on its investment. Another party, $p_{2}$, provides a budget for the project and has veto power. It usually acts as a middle ground between different parties. There exists a group of beneficiary parties, $P_{\text{benefit}}\in P$, whose interests can align with $p_{1}$ in multiple issues, but they want to negotiate better deals. Some parties $P_{\text{const}}\in P$ (e.g., activists, environmentalists) would like to impose more constraints on the project, which usually contradicts $p_{1}$’s interests. Other parties, $P_{\text{oppose}}\in P$, can have mostly opposing interests to $p_{1}$ as the project may have a frictional effect on their operations, living conditions, etc.
@@
Issues. Parties negotiate over 5 issues $I=\{A,B,...,E\}$ that are relevant to the project (e.g., funding, location, revenue, etc.). Each issue has 3-5 sub-options, e.g., $A=\{a_{1},a_{2},...,a_{n}\}$. A deal, $\pi\in\Pi$ where $\Pi$ is the set of all deal combinations, consists of one sub-option per issue, $\pi=[a_{k}\in A,b_{l}\in B,c_{m}\in C,d_{n}\in D,e_{o}\in E]$. In our case, the total number of possible deals $|\Pi|$ is 720. The issues and sub-options can represent a range over a quantity in dispute (e.g., size of the project, allocated fund), or they can take a more discrete form (e.g., different locations). To denote that party $p_{i}$ suggested a deal at a point $t$ during the game, we use the notation $\pi_{p_{i}}^{(t)}$.
@@
Scoring. Each party $p_{i}$ has its scoring system $S_{p_{i}}$ for the sub-options, representing and proportional to the value it assigns to them. The sub-options have semantic connections to the parties’ goals (e.g., will lead to maximizing or minimizing its profit return, etc.). The priority of issues (e.g., $\text{max}(S_{p_{i}}(a_{1}),S_{p_{i}}(a_{2}),...,S_{p_{i}}(a_{n}))$ ) differ between parties, also aligning with the parties’ goals. Some parties might be completely neutral on some issues (indicated by a score of 0). For a party $p_{i}$, its score of a deal (suggested by $p_{j}\in P$) is the sum of its scores of this deal’s sub-options, i.e., $S_{p_{i}}(\pi_{p_{j}}^{(t)})=S_{p_{i}}(a_{k})+S_{p_{i}}(b_{l})+S_{p_{i}}(c_{m})+S_{p_{i}}(d_{n})+S_{p_{i}}(e_{o})$, with a maximum of 100.
@@
Feasible solutions. Each party has a minimum threshold for acceptance; in negotiation terms, this is known as “Best Alternative To a Negotiated Agreement” (BATNA). A deal is feasible if it exceeds the thresholds of at least 5 parties, which must include the project’s proposer and the veto party, $p_{1}$ and $p_{2}$. These factors restrict the set of feasible deals $\Pi_{\text{pass}}\in\Pi$ and can quantify the success in reaching an agreement. They also control the game’s difficulty by increasing/decreasing the size of the feasible set $|\Pi_{\text{pass}}|$, which allows the instantiation of new games, as we show in our experiments.
Newly created games. LLMs have been used to create quests in role-play games (Värtinen et al., 2022) or to automatically create test examples (Gandhi et al., 2023a). Thus, we use LLMs as a seed to instantiate new negotiation games by creating the background story, the parties, the issues, and the goals and preferences of each party (see Appendix E for the prompt we provided to Bing Chat and an example of a new game). We specify that parties should include a proposer, a resource manager, a beneficiary, opposing parties, etc., and issues should represent competing interests of parties over, e.g., a shared resource. This can be thought of as creating a game template that the LLM can populate (Gandhi et al., 2023a), which enables a flexible and LLM-assisted creation of new diverse games. We manually curated these games by rewriting them to follow the first base game’s description style and by changing some of the preferences and sub-options to ensure logical consistency. We often submitted follow-up instructions to the model to rewrite the whole game conditioned on changes that we specified. We then assigned numerical scores for the sub-options. We tuned the scores and thresholds to reach a comparable number of feasible deals $|\Pi_{\text{pass}}|$ compared to the base game (55 to 57 possible deals with at least a 5-way agreement). We provide more discussion on comparing the difficulty of the new games vs. the base one later in our experiments.
Game variants. There are two variants of the game, which we call “All in” and “One out”. In “All in”, all parties are interested in establishing a deal that is higher than their minimum thresholds (with a varying degree of greediness vs. cooperativeness). In “One out”, one party has an incentive to sabotage the negotiation (it will receive a higher score if the negotiation fails). We set this party to be $\in P_{\text{const}}$. Under “All in”, we study sub-variants, where one agent is prompted to be greedy while the others are cooperative, and another where all agents are cooperative (baseline variant).
4 LLMs Playing the Game: Setup and Prompting Strategies
In this section, we present our setup and prompting strategies to solve the game, shown in Figure 2.
4.1 Game Setup
@@
Initial prompts. Each agent $p_{i}$ is characterized via an initial prompt (see Appendix E) that consists of 1) shared information about the project, the parties involved, and the issues descriptions, 2) confidential information about the scores of this particular agent $S_{p_{i}}$ and its minimum threshold, and 3) general instructions explaining the game rules (e.g., not disclosing scores). To make models more likely to associate scores with goals, the initial prompts mention how scores correlate with goals and give 1-2 examples of how other agents’ scores can differ according to their goals.
@@
Rounds. $p_{1}$ starts the negotiation by suggesting its ideal deal. Then, the game continues for $r$ rounds; in each, one agent is randomly selected and prompted with the initial prompt, a history of the most recent $n$ interactions, and instructions that guide the negotiation (more details in the following subsection). Agents should either support previous deals or propose new ones.
@@
End of negotiation. After all rounds, the project’s proposer $p_{1}$ is prompted with instructions to propose a final official deal ($\pi_{p_{1}}^{(r+1)}$). Similar to rounds’ prompts, these instructions are appended to the initial prompt and the last $n$ interactions.
4.2 Prompting Strategy: Incentives and Structure
We next describe our strategy for prompting agents in terms of 1) the incentives provided to them, and 2) the exact structure enforced in the prompt to decompose the task. The incentives differ according to the game variant (All in and One out) and whether cooperation and greediness are explicitly promoted. The corresponding prompts can be found in Appendix F.
Figure 2: Game setup.
Incentives: All in - all cooperative. This is our baseline game variant, in which all agents are instructed that any deal higher than their minimum threshold is preferable to them than no deal. At each round, they are also instructed to aim for a balanced agreement, show flexibility, consider other parties’ interests and accommodate them, and propose deals that are likely to lead to an agreement.
Incentives: All in - one greedy. In this game variant, we introduce a greedy agent and keep the others cooperative. The greedy agent is instructed to maximize its own score and benefits as much as possible and compromise only for issues that do not matter or matter the least to it. However, it is still incentivized to reach a deal; the instructions state to “be careful not to completely ruin the deal because you still want to have a deal that is higher than your BATNA (your minimum threshold).”

### v2

# Cooperation, Competition, and Maliciousness: LLM-Stakeholders Interactive Negotiation

@@ Abstract
There is an growing interest in using Large Language Models (LLMs) in multi-agent systems to tackle interactive real-world tasks that require effective collaboration and assessing complex situations. Yet, we still have a limited understanding of LLMs’ communication and decision-making abilities in multi-agent setups. The fundamental task of negotiation spans many key features of communication, such as cooperation, competition, and manipulation potentials. Thus, we propose using scorable negotiation to evaluate LLMs. We create a testbed of complex multi-agent, multi-issue, and semantically rich negotiation games. To reach an agreement, agents must have strong arithmetic, inference, exploration, and planning capabilities while integrating them in a dynamic and multi-turn setup. We propose multiple metrics to rigorously quantify agents’ performance and alignment with the assigned role. We provide procedures to create new games and increase games’ difficulty to have an evolving benchmark. Importantly, we evaluate critical safety aspects such as the interaction dynamics between agents influenced by greedy and adversarial players. Our benchmark is highly challenging; GPT-3.5 and small models mostly fail, and GPT-4 and SoTA large models (e.g., Llama-3 70b) still underperform.
@@ 1 Introduction
Large Language Models (LLMs) [5, 29] are used in tasks beyond traditional NLP, such as using tools [34, 19, 46] or solving reasoning problems [37, 43]. They are adopted in many real-world applications [27, 22, 23] that require multi-turn interactions and adaptation to external sources and interfaces [27]. Multi-agent LLM frameworks are envisioned to be a key design pattern for future autonomous systems [26]. However, LLMs are not explicitly trained for these tasks. Given this contrast, we need new evaluation frameworks to assess models in complex communication settings.
Complex communication involved in, e.g., satisfying customers, agreeing on contracts, and high-stake decisions, such as authorizing loans, requires prolonged deliberation. We use crucial skills such as strategic planning, competition, cooperation, balancing between multiple objectives, and awareness of cooperation barriers such as manipulation and deception. This should ideally apply to AI and LLM agents, which are increasingly relied on as personal [25, 28] and negotiation assistants [11, 20, 30, 10]. A future where AI assistants communicate on behalf of different entities seems plausible. This raises the concern of models being exploited by rogue parties to pursue unaltruistic or manipulative goals.
@@
Figure 1: Left: Parties negotiate over 5 issues with different sub-options. Each party has its own secret scores, issue priorities, and a minimum threshold for acceptance. Right: Parties ideally reach a common ground by adjusting their optimum deal. This is visible in the graph; over rounds, the leading agent $p_{1}$ proposes deals that reduce its own score but increase all agents’ collective score.
As negotiation is integral to these scenarios [14] and thus for advancing AI agentic design, we propose scorable negotiation games, with complex cooperation and competition between multiple parties, as a multi-step dynamic benchmark for LLMs. In these games, agents ideally assess the value of deals w.r.t. their own goals, have a representation of others’ goals, weigh different options, and finally find common grounds. These sub-tasks require substantial arithmetic and strategic reasoning under only partial observations. They also span commonsense reasoning [40, 32] and Theory-of-Mind (ToM) capabilities [35, 33]. Such skills are required in many applications to rank and propose solutions, e.g., to answer “find the cheapest, shortest flight with a reputable airline that will not lose my luggage".
@@
We first use a role-play exercise commonly used for teaching negotiation [38], which consists of multiple parties and issues (see Figure 1). Parties have their real-world-inspired goals correlated with their individual secret scores for issues. They also have a minimum threshold for agreement. The priorities vary between parties, creating a non-zero-sum game with potential for cooperation and competition. The scores and thresholds control the set of feasible solutions, providing a way to quantify performance. We use an LLM as a seed to design 3 completely new and diverse games from scratch. We easily instantiate new games with different difficulty levels by changing scores and thresholds. These factors make our benchmark highly evolving to test future more powerful models.
We design a baseline framework, via prompting, that systematically breaks down the task into intermediate ones, revealing essential insights about the most needed capabilities. Our findings show that GPT-4 [29] (the best-evaluated model) still underperforms when increasing games’ difficulty. Furthermore, GPT-4 agents can get higher rewards compared to GPT-3.5 [5] ones when assigned the same role in a mixed population simulation, hinting at potential fairness and disparity considerations when users use models with varying capabilities as assistants. Some open-source models (Llama2/3 70b [41, 21] and Mixtral [13]) outperform GPT-3.5 and the latest version of Gemini [3].
Moreover, our complex environment enables us to study agents’ dynamics in unbalanced and adversarial setups, a critical aspect of autonomous agents. We show that agents can be steered toward greediness or manipulation, altering other compromising agents’ behaviors, which may reward the greedy agent’s demands more highly. The adversarial agent may also create a coalition against a target agent, etc. These attacks are broadly useful for AI safety research to study AI manipulation and deception [31], alignment of multi-agent systems, and actions driven by an assigned persona [2, 36].
In summary, our work provides several complex and interactive negotiation games as an evolving benchmark to test LLMs’ capabilities, the potential for manipulation, and future robustification. To foster future research, we will release our toolkit of diverse games, code platform, and transcripts.
@@ 2 Game Description
Games consist of 6 parties, $P=\{p_{1},p_{2},...,p_{6}\}$, and 5 issues $I=\{A,B,...,E\}$ with dynamics outlined below. All notations and prompts are in Appendices A and I.
@@
Parties. An entity $p_{1}$ proposes a project (e.g., an airport) that it will manage and invest in and wants to increase the return on its investment. Another party, $p_{2}$, provides a budget for the project and has veto power. It usually acts as a middle ground between different parties. There exists a group of beneficiary parties, $P_{\text{benefit}}\in P$, whose interests can align with $p_{1}$ in multiple issues, but they want to negotiate better deals. Some parties $P_{\text{const}}\in P$ (e.g., environmentalists) would like to impose more constraints on the project, which usually contradicts $p_{1}$’s interests. Other parties, $P_{\text{oppose}}\in P$, have opposing interests to $p_{1}$ as the project may affect their operations, living conditions, etc.
@@
Issues. Parties negotiate over 5 issues $I=\{A,B,...,E\}$ related to the project (e.g., funding). Each issue has 3-5 sub-options, e.g., $A=\{a_{1},a_{2},...,a_{n}\}$. A deal, $\pi\in\Pi$ where $\Pi$ is the set of all deal combinations, consists of one sub-option per issue, $\pi=[a_{k}\in A,b_{l}\in B,c_{m}\in C,d_{n}\in D,e_{o}\in E]$. The total number of possible deals $|\Pi|$ is 720. The sub-options take the form of a range over a quantity in dispute (e.g., project size, revenue, etc.) or a discrete form with less apparent compromise (e.g., different locations). To denote that party $p_{i}$ suggested a deal at a time $t$, we use the notation $\pi_{p_{i}}^{(t)}$.
@@
Scoring. Each party has its own scoring system $S_{p_{i}}$ for the sub-options, which has a semantic connection to the parties’ goals (e.g., will increase or decrease its profit return). The priority of issues (e.g., $\text{max}(S_{p_{i}}(a_{1}),S_{p_{i}}(a_{2}),...,S_{p_{i}}(a_{n}))$ ) differ between parties. Some parties can be completely neutral on some issues (indicated by a score of 0). These factors result in a non-zero-sum game and control the cooperation and competition between parties. For a party $p_{i}$, its score of a deal (suggested by $p_{j}\in P$) is the sum of its scores of this deal’s sub-options, i.e., $S_{p_{i}}(\pi_{p_{j}}^{(t)})=S_{p_{i}}(a_{k})+S_{p_{i}}(b_{l})+S_{p_{i}}(c_{m})+S_{p_{i}}(d_{n})+S_{p_{i}}(e_{o})$, with a maximum of 100.
@@
Feasible solutions. Each party $p_{i}$ has a minimum threshold $\tau_{p_{i}}$ for acceptance. A deal is feasible if it exceeds the thresholds of at least 5 parties, which must include $p_{1}$ and $p_{2}$. These factors restrict the set of feasible deals $\Pi_{\text{pass}}\in\Pi$, quantify the success in reaching an agreement, and control the game’s difficulty by altering the size of the feasible set $|\Pi_{\text{pass}}|$, which allows instantiating new games.
New games. The base game is adapted, with our own descriptions, from a negotiation exercise [38, 39]. Moreover, we use LLMs to create new games by creating the background story, the parties, the issues, and the goals and preferences of each party, from scratch; the base game is not given to the model as in-context information. We only specify that parties should include a proposer, a resource manager, a beneficiary, opposing parties, etc., and issues should represent competing interests of parties. We manually curated the games to ensure logical consistency, and we assigned numerical scores to reach a comparable number of feasible deals compared to the base game ($\sim$55 deals).
3 LLMs Playing the Game
We here present agents’ interaction protocol, the different variants of the game, and our prompting solution framework. Our setup is in Figure 2. Algorithm and prompts are in Appendices A and J.
Figure 2: Interaction protocol.
3.1 Agents’ Interaction Protocol
@@
Initial prompts. Each agent $p_{i}$ is characterized via an initial prompt that consists of 1) shared information about the project, the parties involved, and the issues’ descriptions, 2) confidential information about the scores of this particular agent $S_{p_{i}}$ and its minimum threshold $\tau_{p_{i}}$, and 3) general instructions explaining the game rules (e.g., not disclosing scores). The initial prompts mention how scores correlate with goals and give 1-2 examples of how other agents’ scores can differ according to their goals.
@@
Rounds. $p_{1}$ starts the negotiation by suggesting its ideal deal. The game then continues for $R$ rounds; in each, one agent is prompted with the initial prompt, a history of the most recent $n$ interactions, and rounds’ instructions that guide the negotiation (more details in the following). Agents should either support previous deals or propose new ones. The input context and output of agent $p_{i}$ at time $t$ are:
$O_{p_{i}}^{(t)}=\text{LM}(C_{p_{i}}^{(0)},H^{(-n)},C_{p_{i}}^{(t)}),$ (1)
$H^{(-n)}$ is the most recent $n$ public answers, $C_{p_{i}}^{(0)}$ is the initial prompt, and $C_{p_{i}}^{(t)}$ is the rounds’ prompt.
@@
End of negotiation. After $R$ rounds, the project proposer $p_{1}$ is prompted with instructions to propose a final official deal ($\pi_{p_{1}}^{(R+1)}$). Similar to eqn. 1, these instructions are appended to the initial prompt and the last $n$ interactions. This final deal determines whether an agreement has been reached. The achieved utility of each party becomes:
$U_{p_{i}}=\left\{\begin{array}[]{ll}S_{p_{i}}(\pi_{p_{1}}^{(R+1)})&\text{if }\pi_{p_{1}}^{(R+1)}\in\Pi_{\text{pass}}\\ \text{BATNA}&\text{otherwise,}\\ \end{array}\right.$ (2)
where BATNA is Best Alternative To a Negotiated Agreement, which is usually the threshold $\tau_{p_{i}}$ but may differ depending on the game variants outlined next.
3.2 Compromising, Greedy, and Adversarial Games
The agents’ scores entail different levels of cooperation and competition. For example, the game will be more competitive if all parties equally prioritize the same issue with very opposing interests. In addition to that, we further evaluate how agents’ actions can be explicitly modulated to promote compromise, greediness, or maliciousness.
Compromising game. Here, all agents are instructed that any deal likely to lead to an agreement and higher than their minimum threshold is preferable to no deal; i.e., the BATNA of agents in eqn. 2 is their minimum threshold. Specifically, the optimization problem an agent $p_{i}$ performs is modeled as:
$f(\pi)=w_{p_{i}}S_{p_{i}}(\pi)+\sum_{p_{j}\in P\backslash\{p_{i}\}}{w_{p_{j}}S_{p_{j}}^{*}(\pi)}$ (3)
$\pi_{p_{i}}^{(t)}\coloneqq\underset{{\pi\in\{S_{p_{i}}(\pi)>\tau_{p_{i}}\}}}{\arg\max}f(\pi);$ (4)
$p_{i}$ cannot observe the scores of another agent $p_{j}$. Therefore, $S^{*}$ is $p_{i}$’s estimate. $w_{p_{i}}$ and $w_{p_{j}}$ are weights assigned to the agent’s own score vs. $p_{j}$’s. The agent may prioritize some agents (e.g., veto parties) over others. In the compromising game, the agent is not particularly prioritizing its own score over others; $w_{p_{i}}\leq\text{min}(\{w_{p_{j}}|\;p_{j}\in P\backslash\{p_{i}\}\})$.
Greedy game. When agents interact in the real world with other agents or humans, they might face non-collaborative or even exploitative players. Thus, we introduce one or more greedy agents and keep the others compromising. The greedy agents are instructed to maximize their own score and benefits as much as possible while still aiming for an agreement; i.e., the BATNA is still the minimum threshold. The optimization objective is similar to eqn. 3, but with $w_{p_{i}}\gg\text{max}(\{w_{p_{j}}|\;p_{j}\in P\backslash\{p_{i}\}\})$.